macgirl234:

Compare the following functions: f(x) = 2 sin(2x − π) + 2 g(x) graph of a quadratic with points at 1, 2 and 3, negative 2 and 5, 2 h(x) x y −2 10 −1 7 0 5 1 3 2 5 3 7 4 10 Which function has the smallest minimum?

Hero:

@macgirl234 have you graphed f yet?

macgirl234:

No. I have no idea where to begin with any of this

Hero:

Okay, no problem

macgirl234:

Can you help me?

Hero:

Yes, so for each function, you have inputs and outputs. The output values are the actual values of a function for its corresponding input. For example, The value of h is 5 when x equal 0.

Hero:

What you want to do is make a table of values for each function, then determine which function has the smallest minimum value.

macgirl234:

I still don't understand...

Hero:

Hang on a minute

macgirl234:

ok

Hero:

Okay, sorry for taking so long

Hero:

First here is a graph of f(x): https://www.desmos.com/calculator/clkwu15jna

Hero:

And a table of values for f(x) 1497833763-594718cd57b07fc89e303000-Tableoff(x).png

Hero:

The rightmost column represents the values for f which includes minimum values, maximum values and everything in between.

Hero:

Plotting the table for g(x) is a bit more involved because we have to actually find the equation for it. They give us three points on its graph (1,2), (3,-2) and (5,2). We have to take those three points and insert them in to \(g(x) = ax^2 + bx + c\) like so 2 = a(1)^2 + b(1) + c -2 = a(3)^2 + b(3) + c 2 = a(5)^2 + b(5) + c When we simplify that we have 2 = a + b + c -2 = 9a + 3b + c 2 = 25a + 5b + c We have to solve for a, b, and c to get the coefficients of g(x).

Hero:

Once we get those coefficients, we can then graph g(x) and find its minimum value.

Hero:

(1st) 2 = a + b + c (2nd) -2 = 9a + 3b + c (3rd) 2 = 25a + 5b + c Subtracting the 1st equation from the 2nd we get: -4 = 8a + 2b Subtracting the 2nd equation from the 3rd we get: 4 = 16a + 2b In other words, we end up with the following system: -4 = 8a + 2b 4 = 16a + 2b Which simplifes to -2 = 4a + b 2 = 8a + b Subtracting the top equation from the bottom we get: 4 = 4a So a = 1 Since b = 2 - 8a, we can find b: b = 2 - 8(1) = 2 - 8 = -6 So b = -6 Since c = 2 - a - b, we can find c as follows: c = 2 - 1 -(-6) = 2 + 6 - 1 = 8 - 1 = 7 So a = 1, b = -6, and c = 7, therefore \(g(x) = x^2 -6x + 7\)

Hero:

Here's the graph of g(x): https://www.desmos.com/calculator/m3xtksexsz

Hero:

And it is clear that the lowest value of g(x) is -2 1497835062-594718cd57b07fc89e303000-Graphofg(x).JPG

Hero:

Here's the graph of h which has a minimum value of 3 https://www.desmos.com/calculator/i6icbyu96p

Hero:

So @macgirl234 given the above information, which function has the lowest minimum?